OpenStudy (anonymous):

Can you help? A medal will be given! :3 Part A: Use the properties of exponents to explain why 64 raised to the power of 1 over 2 is called the square root of 64. Part B: The length of a rectangle is 7 units and its width is square root of 7 unit. Is the area of the rectangle rational or irrational? Justify your answer.

OpenStudy (anonymous):

Part A \(\Huge 64^{\frac{1}{2}}=(8×8)^{\frac{1}{2}} \) now for the properties of exponents we chose the have numbers of them so \(\Huge 64^{\frac{1}{2}}=(8×8)^{\frac{1}{2}}=8=\sqrt{64} \)

OpenStudy (anonymous):

part B  \(\Huge\text{ the area = length × width}\) \(\Huge\text{ the area} = 7 ×\sqrt {7}=\color{red}{7\sqrt7}=7^{\frac{3}{2}}\) so its irrational , u cant write \(7^{\frac{3}{2}}\) as \(\frac{a}{b}\) formula

OpenStudy (anonymous):

got it ? if not , just let me know :D

OpenStudy (anonymous):

I got it, thank you so much :3

OpenStudy (anonymous):

np ur wlc ^^

OpenStudy (anonymous):

wait, for part A, why is it a square root?

OpenStudy (anonymous):

becuz , u chose one of two ^^

OpenStudy (anonymous):

OOOOh, ok, thanjs